# Fast and Globally Consistent Normal Orientation based on the Winding Number Normal Consistency

### v1

@@ 5.1.1 Implementation Details
In our implementation, the input point cloud is normalized to fit into the cube $[-1,1]^{3}$ with a margin of 1/$1/1$ 11 from the cubes boundaries. The smoothing widths are set as $w_{1}=0.002$ and $w_{2}=0.016$ unless otherwise stated. In Algorithm 4, the maximum tree depth is $D=15$, and the threshold for using an representative source is $c=2$. We run 40 iterations for each model.
5.1.2 Baselines
@@
We choose recent state-of-the-art methods dipole propagation [34] iterative Poisson Surface Reconstruction (iPSR) [19], Parametric Gauss Reconstruction (PGR) [31] and GCNO [44] as the baselines. We use their default parameters for our experiments unless otherwise mentioned. For Dipole [34], we use orient_large for point clouds with more than 50000 points, and orient_pointcloud otherwise. All experiments are done on a desktop computer with an Intel i9-10900X CPU (3.70GHz, 10 cores) and an NVIDIA RTX 3090 GPU.
5.1.3 Datasets
@@
Following previous work, we use datasets that are commonly used in surface reconstruction and normal orientation, including ABC [26], Thingi10k [47], the Stanford scanning repository [8, 28, 39, 13], and other datasets from Lu et al. [32], Huang et al. [20, 21], Erler et al. [12], Metzer et al. [34], Lin et al. [31], Xu et al. [44].
5.1.4 Metrics
@@
${\rm AE}_{\rm pcd}=\frac{1}{N}\sum_{i=1}^{N}(1-n_{i}^{\rm gt}\cdot n_{i}^{\rm recon})/2.$ (19)
We also follow Xu et al. [44] to use the percentage of correctly oriented normals as a metric:
$P_{\rm co}=\#\{i:n_{i}^{\rm recon}\cdot n_{i}^{\rm gt}>0\}/N.$ (20)
Another common practice is to measure the accuracy of normals using their reconstructed meshes. For fairness, we use Screened Poisson Surface Reconstruction (SPSR) [23] for all methods. We set the maximum tree depth in SPSR to 10 and keep other parameters as default. We use the mesh Chamfer distance (CD) and mesh normal angular error (${\rm AE}_{\rm mesh}$). Here, ${\rm CD}$ for two meshes $M_{1},M_{2}$ is defined as
${\rm CD}=\frac{1}{2}\Big{(}\frac{1}{N}\sum_{i=1}^{N}\|x_{i}-\hat{x}_{i}\|+\frac{1}{N}\sum_{i=1}^{N}\|y_{i}-\hat{y}_{i}\|\Big{)},$ (21)
@@
GT normalsOursiPSRDipolePGRGCNO00.5
Figure 5: Normal orientation results of different methods. We show the point clouds colored by their normals and by per-point angular error values (Eq. (19) without averaging across points). We also show the histograms of angular errors ($y$-axis in log-scale).
GT mesh and normalsOursiPSRDipolePGRGCNO
Figure 6: Normal orientation and surface reconstruction results of different methods. We show reconstructed meshes in plain colors and point clouds colored by their normals.
Table 2: Quantitative comparison results of different methods evaluated with uniform samples. The best results are marked in bold.
Method Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$ Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$
(×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%) (×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%)
GCNO 38.1440 11.3096 17.7151 88.4400 29.7463 12.5233 9.3252 98.3400
PGR 13.1959 11.5137 18.0785 86.8780 XYZRGB 13.5690 12.9512 21.8533 84.3820
Dipole Crab 15.1305 12.7237 12.2426 88.6625 Statuette 14.9317 13.5039 15.6297 86.2206
iPSR 2.0727 1.6770 1.9529 99.9356 2.7776 2.8389 3.4912 99.8594
Ours 1.4556 1.3515 0.8401 99.9981 2.0174 2.4180 1.7715 99.9781
GCNO 15.2068 7.2695 7.4188 99.2200 24.6961 6.7809 5.2512 99.2000
PGR 4.6810 4.3906 13.5943 93.2160 Statue of 7.4997 5.9834 11.5123 93.4080
Dipole Liberty 4.8793 4.0816 4.6036 96.2181 Delaunay 145.0533 30.4131 39.6944 60.5975
iPSR 1.5647 1.2212 1.3644 99.9963 1.5918 1.0918 1.0650 99.9963
Ours 1.0802 1.0209 0.4961 99.9975 1.1420 0.9144 0.4398 99.9994
GCNO 65.8343 44.8330 47.0361 53.9437 18.3662 8.8840 8.7805 97.6000
PGR Trash 6.4437 0.6648 2.1229 99.6740 7.9294 8.9578 19.2768 88.4840
Dipole Can 34.4190 25.5114 22.1820 77.9050 Dragon 4.0927 5.7576 5.3654 95.5606
iPSR 23.9465 14.9586 13.9820 86.4475 1.9667 5.1030 5.0781 96.3306
Ours 1.2967 0.3240 0.1370 100.0000 1.3937 4.9502 4.1447 96.3306
Table 3: Quantitative comparison results of different methods on a large-scale benchmark dataset [21]. Each metric is an average over all 1387 models used for this evaluation. The best results are in bold.
@@
Figure 7: Qualitative comparisons for models with thin structures and sharp edges. Each model in the first two rows has 100,000 points, and each in the last two rows has 10,000 points.
@@
Figure 8: Reconstruction results of two high-genus surfaces. We additionally show the genus $g$ of the GT models and the reconstructed meshes.
5.2 Evaluation of Normal Quality
5.2.1 Accuracy
We use uniform samples without noise to measure the accuracy of different methods. For each mesh, we use sample_surface_even in Trimesh [9] to obtain uniform samples. Each model is sampled with 160,000 points. Note that PGR [31] and GCNO [44] have prohibitively high memory and time consumption, respectively. Hence, we downsample the input point cloud to 50,000 points for PGR and 5,000 points for GCNO. A more fair comparison is provided separately in Sec. 5.2.6.
Fig. 5 shows the oriented point clouds, their angular error maps and angular error distribution. It can be seen that dipole propagation [34] has a high propability of incorrectly orienting normals in a large area. For other methods, while they can generally achieve a globally consistent orientation, their estimated normals have larger errors compared to our results. Fig. 6 shows reconstructed meshes using normals estimated by different methods. Note that almost all baselines cannot handle convex thin structures (see the interior of the trash can model). The quantitative results for these six models are reported in Table 2. Our method achieves the best overall performance.
To further validate the accuracy of our estimated normals, we conduct a large-scale study on a new surface reconstruction benchmark dataset [21]. Since we found the point clouds provided by the Huang et al. [21] contain noise (in both point positions and normals) and are unfit for quantitative evaluation, we use their provided meshes and sample uniform points using Trimesh [9]. Furthermore, some meshes provided by this dataset are non-manifold with self-intersecting components, which are not suitable for either our method or the baselines. Hence, we remove all non-manifold data and use the remaining 1387 models for this large-scale evaluation. Each model is sampled with 160,000 points. Note that PGR [31] and GCNO [44] are too computationally expensive and are thus excluded from this comparison. Table 3 reports the metrics averaged across all 1387 models. Our methods outperforms iPSR [19] and Dipole [34] in a statistical sense.
@@ 5.2.1 Accuracy
Figure 9: Normal estimation and reconstruction results of sparse point clouds and 3D sketches from Huang et al. [20]. Oriented points are rendered as point sprites, and we additionally render normals for 3D sketches.
5.2.2 Thin Structures and Sharp Edges
@@
5.2.3 High-genus Surfaces
We experiment with two high-genus surfaces shown in Fig. 8. Due to their large surface areas, we sample 200,000 for each model to ensure the point clouds can depict the underlying geometric structure. We emprically found that our method and iPSR require more iterations to converge, and thus set the number of iterations for both methods to 300. In Fig. 8 we show the reconstruction results as well as the genus of each mesh. Our method can more faithfully reconstruct topological structures than other methods.
5.2.4 Sparse Points and 3D Sketches
Sparse points can be another challenge due to low sampling rates. Additionally, Huang et al. [20] studied an extreme case of sparse inputs: 3D sketches, where the points are not only sparse but also highly uniform. In Fig. 9 we show four models from Huang et al. [20]. The two uniform sparse point clouds each have 500 points, and the two sketches each have 1,000 points. Our method can obtain plausible normals for both types of point clouds.
[…]
Figure 10: Reconstruction (in yellow) and normal estimation results (in green) for a chair model with different noise levels. The normals are rendered as point sprites to indicate their orientation.
Input pointsGT meshOursiPSRDipolePGRGCNO
Figure 11: Reconstruction results for four noisy point clouds. Each model is obtained by adding $\sigma=0.5\%$ Gaussian noise to uniform samples.
Table 4: Quantitative comparisons of estimating normals for noisy models exhibited in Fig. 10 and Fig. 11. Since it does not make sense to talk about “GT normals” of noisy points, we measure only mesh errors. The best results are marked in bold.
@@ 5.2.4 Sparse Points and 3D Sketches
GCNO 8.3960 9.8521 3.7648 5.4055 4.2233 11.8817 3.0652 3.2314
PGR 7.3970 10.3166 3.2521 4.8484 4.4460 18.8665 3.0096 8.6615
Dipole Cup 5.7139 23.4262 Flower 6.4231 29.9646 Buddha 5.1984 32.2373 Vase 20.2586 47.9005
@@
5.2.5 Noise Resilience
We study the noise resilience of different methods by adding Gaussian noise to clean uniform samples. We use three noise levels $\sigma=0.25\%,0.5\%,1\%$ (w.r.t. the bounding box diagonal) as the standard deviation for Gaussian noise. Following previous settings, each model is sampled with 160,000 points, which are downsampled to 50,000 for PGR [31] and 5,000 for GCNO [44]. We noticed that default settings of several methods do not work well with noisy inputs, and therefore we adjust the parameters to obtain optimal results. For our method, we set $w_{1}=0.02,0.03,0.04$ and $w_{2}=4w_{1}$ for the three noise levels. For iPSR [19], we set the point weight to 0.5. For PGR [31], we use their recommended settings for noisy data. Note that Dipole [34] and GCNO [44] do not have noise-adapting options or parameters, and we keep their default settings. We also change the point weight to 0.5 for the final reconstruction step using SPSR [23].
Fig. 10 shows the reconstructed meshes (in yellow) and the estimated normals (in green) for different noise levels. For dipole propagation [34], even without noise it cannot infer normals with globally consistent orientation, and as noise level increases its normals become completely misoriented. Other methods can infer a roughly correct orientation but are notably worse at preserving the geometric structure, especically at high noise levels. Fig. 11 exhibits four more noisy cases (all models are sampled with noise level $\sigma=0.5\%$). Compared with other methods, ours achieves the best balance between preserving the overall topological structure and surface details. Table 4 reports the quantitative results for the models in Fig. 10 and Fig. 11.
5.2.6 Extended Comparisons with GCNO and PGR
In our previous experiments, we have used downsampled point clouds for GCNO [44] and PGR [31] due to their high complexity. To make the comparison with PGR and GCNO more comprehensive, we rerun our method for all models used in Table 2 and Table 4 with the same number of samples as PGR and GCNO. The new metrics are reported in Table 5 and Table 6. Note that our method still outperforms these two baselines.
Fig. 12 shows the qualitative comparisons between our method, PGR and GCNO using the same number of samples. Furthermore, the authors of PGR [31] suggested that the normals solved by PGR are not really compatible with SPSR [23], and better reconstructions can only be obtained by plugging these normals back into the winding number field (WNF). While we have previously used SPSR as the reconstruction for all methods for fairness, we also test PGR with WNF-based reconstruction as in Lin et al. [31]. In Fig. 12, we observe that our method still outperforms PGR despite the final reconstruction method.
5.2.7 Real-world Scans
Different from synthetic data, real-world scans generally suffers from several kinds of imperfections, such non-uniform point distribution, scanner noise, outlier points, multi-view misalignment, occluded missing regions, etc. Furthermore, scanner data are usually large in scale, having millions of points. To evaluate the practicability of our method on scanned data, we experiment with three raw range scans (Armadillo, Bunny and Dragon) from the Stanford scanning respository [8, 28, 39, 13] and one scan (Lady) from Lu et al. [32]. Note that some raw scans contain background points as well, and we manually set a bounding box to filter out background points.
Table 5: Quantitative comparison results between our method, PGR and GCNO using the same number of samples. For comparisons with PGR and GCNO we use 50,000 points and 5,000 points for each model, respectively.
Method Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$ Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$
(×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%) (×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%)
GCNO 38.1440 11.3096 17.7151 88.4400 29.7463 12.5233 9.3252 98.3400
Ours Crab 20.7035 7.3027 5.7689 98.7800 XYZRGB 26.5084 12.2513 9.0382 98.0600
PGR 13.1959 11.5137 18.0785 86.8780 Statuette 13.5690 12.9512 21.8533 84.3820
Ours 3.4914 2.6714 1.8335 99.9300 5.0047 4.8284 3.4662 99.8560
GCNO 15.2068 7.2695 7.4188 99.2200 24.6961 6.7809 5.2512 99.2000
Ours Liberty 12.5331 6.8335 5.9479 99.6800 Statue of 20.6545 6.3045 4.2547 99.4400
PGR 4.6810 4.3906 13.5943 93.2160 Delaunay 7.4997 5.9834 11.5123 93.4080
Ours 3.4957 2.6623 1.2681 99.9880 3.3201 2.1595 1.0475 99.9920
GCNO 65.8343 44.8330 47.0361 53.9437 18.3662 8.8840 8.7805 97.6000
Ours Trash 102.9334 36.8877 32.5923 70.7071 Dragon 15.2683 8.2227 7.5333 98.0800
PGR Can 6.4437 0.6648 2.1229 99.6740 7.9294 8.9578 19.2768 88.4840
Ours 6.0636 0.7209 0.3859 100.0000 4.3697 5.7403 5.0017 96.3240
Table 6: Quantitative comparisons between our method, PGR and GCNO in the case of noisy inputs (same as the models used in Table. 4). For comparisons with PGR and GCNO we use 50,000 points and 5,000 points for each model, respectively. The reported metrics are for the models shown in Fig. 10 and Fig. 11. Since it does not make sense to talk about “GT normals” of noisy points, we measure only mesh errors.
GCNO 4.1716 0.5879 2.8376 0.6831 3.9593 0.5879 7.7393 0.6831
Ours 0 2.5960 0.5502 0.25% 2.5552 0.6374 0.5% 3.3851 0.5502 1% 5.3490 0.6374
PGR 1.5540 0.2830 1.5270 0.5409 9.3366 0.2830 6.6418 0.5409
Ours 0.3961 0.1694 1.0456 0.3223 4.9580 0.1694 3.4486 0.3223
@@ 5.2.7 Real-world Scans
Ours Cup 5.0091 5.2786 Flower 3.3762 4.8993 Buddha 3.8669 10.8446 Vase 2.6241 2.6057
PGR 7.3970 10.3166 3.2521 4.8484 4.4460 18.8665 3.0096 8.6615
Ours 2.5295 2.2244 1.9771 3.8066 2.3528 8.2716 1.3408 2.3336
GT meshOursPGR normals+SPSRPGR normals+WNFGT meshOursGCNO
Figure 12: We conduct comparisons with GCNO and PGR using the same number of samples. The two models are reconstructed using 9,000 and 50,000 points, respectively. Furthermore, we additionally present the PGR reconstruction result using the winding number field (WNF) based reconstruction.
@@
Figure 13: Reconstruction results of raw range scans from the Stanford scanning repository [8, 28, 39, 13] and Lu et al. [32]. Number of points in each model: 2,374,290 (Armadillo); 362,272 (Bunny); 2,109,047 (Dragon); 530,497 (Lady). Please zoom in to see scanner noise patterns and reconstruction details.

### v2

@@ 5.1.1. Implementation Details
In our implementation, the input point cloud is normalized to fit into the cube $[-1,1]^{3}$ with a margin of 1/$1/1$ 11 from the cube boundaries. The smoothing widths are set as $w_{1}=0.002$ and $w_{2}=0.016$ unless otherwise stated. In Algorithm 4, the maximum tree depth is $D=15$, and the threshold for using a representative source is $c=2$. We run 40 iterations for each model.
5.1.2. Baselines
@@
We choose several recent state-of-the-art methods, dipole propagation (Metzer et al., 2021), iterative Poisson Surface Reconstruction (iPSR) (Hou et al., 2022), Parametric Gauss Reconstruction (PGR) (Lin et al., 2022) and GCNO (Xu et al., 2023), as the baselines. We use their default parameters for our experiments unless otherwise mentioned. For Dipole (Metzer et al., 2021), we use orient_large for point clouds with more than 50,000 points, and orient_pointcloud otherwise. All experiments are done on a desktop computer with an Intel i9-10900X CPU (3.70GHz, 10 cores) and an NVIDIA RTX 3090 GPU.
GTOursiPSRDipoleOursPGROursGCNO00.5
Figure 5. Normal orientation results of different methods. We show the point clouds colored by their normals and by per-point angular error values (Eq. (19) without averaging across points). We also show the histograms of angular errors ($y$-axis in log-scale). Downsampled to 50,000 points. Downsampled to 5,000 points.
[…]
Qualitative results of normal errors.
5.1.3. Datasets
@@
Following previous work, we use datasets that are commonly used in surface reconstruction and normal orientation, including ABC (Koch et al., 2019), Thingi10k (Zhou and Jacobson, 2016), the Stanford scanning repository (Curless and Levoy, 1996; Krishnamurthy and Levoy, 1996; Turk and Levoy, 1994; Gardner et al., 2003), and other datasets from Lu et al. (2018); Huang et al. (2019, 2022); Erler et al. (2020); Metzer et al. (2021); Lin et al. (2022); Xu et al. (2023).
GTOursiPSRDipoleOursPGROursGCNO
Figure 6. Normal orientation and surface reconstruction results of different methods. We show reconstructed meshes in plain colors and point clouds colored by their normals. Downsampled to 50,000 points. Downsampled to 5,000 points.
[…]
Qualitative results of normal orientation and surface reconstructions.
Table 2. Quantitative comparison results of different methods evaluated with uniformly sampled models exhibited in Fig. 5 and Fig. 6. The best results are marked in bold. Downsampled to 50,000 points. Downsampled to 5,000 points.
Method Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$ Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$ Model CD ${\rm AE}_{\rm mesh}$ ${\rm AE}_{\rm pcd}$ $P_{\rm co}$
(×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%) (×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%) (×$\times 10^{-4}$) (×$\times 10^{-2}$) (×$\times 10^{-2}$) (%)
Dipole 15.1305 12.7237 12.2426 88.6625 14.9317 13.5039 15.6297 86.2206 4.8793 4.0816 4.6036 96.2181
iPSR 2.0727 1.6770 1.9529 99.9356 2.7776 2.8389 3.4912 99.8594 1.5647 1.2212 1.3644 99.9963
Ours 1.4556 1.3515 0.8401 99.9981 XYZRGB 2.0174 2.4180 1.7715 99.9781 1.0802 1.0209 0.4961 99.9975
PGR Crab 13.1959 11.5137 18.0785 86.8780 Statuette 13.5690 12.9512 21.8533 84.3820 Liberty 4.6810 4.3906 13.5943 93.2160
Ours 3.4914 2.6714 1.8335 99.9300 5.0047 4.8284 3.4662 99.8560 3.4957 2.6623 1.2681 99.9880
GCNO 38.1440 11.3096 17.7151 88.4400 29.7463 12.5233 9.3252 98.3400 15.2068 7.2695 7.4188 99.2200
Ours 20.7035 7.3027 5.7689 98.7800 26.5084 12.2513 9.0382 98.0600 12.5331 6.8335 5.9479 99.6800
Dipole 145.0533 30.4131 39.6944 60.5975 34.4190 25.5114 22.1820 77.9050 4.0927 5.7576 5.3654 95.5606
iPSR 1.5918 1.0918 1.0650 99.9963 23.9465 14.9586 13.9820 86.4475 1.9667 5.1030 5.0781 96.3306
Ours Statue of 1.1420 0.9144 0.4398 99.9994 Trash 1.2967 0.3240 0.1370 100.0000 1.3937 4.9502 4.1447 96.3306
PGR Delaunay 7.4997 5.9834 11.5123 93.4080 Can 6.4437 0.6648 2.1229 99.6740 Dragon 7.9294 8.9578 19.2768 88.4840
Ours 3.3201 2.1595 1.0475 99.9920 6.0636 0.7209 0.3859 100.0000 4.3697 5.7403 5.0017 96.3240
GCNO 24.6961 6.7809 5.2512 99.2000 65.8343 44.8330 47.0361 53.9437 18.3662 8.8840 8.7805 97.6000
Ours 20.6545 6.3045 4.2547 99.4400 102.9334 36.8877 32.5923 70.7071 15.2683 8.2227 7.5333 98.0800
5.1.4. Metrics
@@
(19) ${\rm AE}_{\rm pcd}=\frac{1}{N}\sum_{i=1}^{N}(1-n_{i}^{\rm gt}\cdot n_{i}^{\rm recon})/2.$
We also follow Xu et al. (2023) to use the percentage of correctly oriented normals as a metric:
(20) $P_{\rm co}=\#\{i:n_{i}^{\rm recon}\cdot n_{i}^{\rm gt}>0\}/N.$
Another common practice is to measure the accuracy of normals using their reconstructed meshes. For fairness, we use Screened Poisson Surface Reconstruction (SPSR) (Kazhdan and Hoppe, 2013) for all methods. We set the maximum tree depth in SPSR to 10 and keep other parameters as default. We use the mesh Chamfer distance (CD) and mesh normal angular error (${\rm AE}_{\rm mesh}$). Here, ${\rm CD}$ for two meshes $M_{1},M_{2}$ is defined as
(21) ${\rm CD}=\frac{1}{2}\Big{(}\frac{1}{N}\sum_{i=1}^{N}\|x_{i}-\hat{x}_{i}\|+\frac{1}{N}\sum_{i=1}^{N}\|y_{i}-\hat{y}_{i}\|\Big{)},$
@@
5.2. Evaluation of Normal Quality
5.2.1. Accuracy
We use uniform samples without noise to measure the accuracy of different methods. For each mesh, we use sample_surface_even in Trimesh (Dawson-Haggerty et al., 2019) to obtain uniform samples. Each model is sampled with 160,000 points. Note that PGR (Lin et al., 2022) and GCNO (Xu et al., 2023) have prohibitively high memory and time consumption, respectively. Hence, we downsample the input point cloud to 50,000 points for PGR and 5,000 points for GCNO. Downsampled experiments are denoted by X (50,000 points) and X (5,000 points), where X is the method name.
Table 3. Quantitative comparison results of different methods on a large-scale benchmark dataset (Huang et al., 2022). Each metric is an average over all 1387 models used for this evaluation. The best results are in bold. Downsampled to 50,000 points.
@@ 5.2.1. Accuracy
Figure 7. Qualitative comparisons for models with thin structures and sharp edges. Each model in the first two rows has 100,000 points, and each in the last two rows has 10,000 points.
@@
Figure 8. Reconstruction results of two high-genus surfaces. We additionally show the genus $g$ of the GT models and the reconstructed meshes.
[…]
Reconstruction results of high-genus surfaces.
@@
Figure 9. Normal estimation and reconstruction results of sparse point clouds and 3D sketches from Huang et al. (2019). Oriented points are rendered as point sprites, and we additionally render normals for 3D sketches.
[…]
Normal estimation and reconstruction results of sparse point clouds and 3D sketches.
Fig. 5 shows the oriented point clouds, their angular error maps and angular error distribution. It can be seen that dipole propagation (Metzer et al., 2021) has a high probability of incorrectly orienting normals in a large area. For other methods, while they can generally achieve a globally consistent orientation, their estimated normals have larger errors compared to our results. Fig. 6 shows reconstructed meshes using normals estimated by different methods. Note that almost all baselines cannot handle convex thin structures (see the interior of the trash can model). The quantitative results for these six models are reported in Table 2. Our method achieves the best overall performance. Note that for the downsampled Trash Can model with 5,000 points, the low sampling rate is insufficient to depict the underlying thin-layer geometry, and thus neither our method nor GCNO could work.
To further validate the accuracy of our estimated normals, we conduct a large-scale study on a new surface reconstruction benchmark dataset (Huang et al., 2022). Since we found the point clouds provided by the Huang et al. (2022) contain noise (in both point positions and normals) and are unfit for quantitative evaluation, we use their provided meshes and sample uniform points using Trimesh (Dawson-Haggerty et al., 2019). Furthermore, some meshes provided by this dataset are non-manifold with self-intersecting components, which are not suitable for either our method or the baselines. Hence, we remove all non-manifold data and use the remaining 1387 models for this large-scale evaluation. Each model is sampled with 160,000 points. Note that PGR (Lin et al., 2022) is memory-intensive, and thus we use inputs downsampled to 50,000 points for this evaluation. Since GCNO (Xu et al., 2023) is too computationally expensive (it would take months even with downsampled points), it is excluded from this comparison. Table 3 reports the metrics averaged across all 1387 models. Our method outperforms the baselines in a statistical sense.
[…]
Figure 10. Reconstruction (in yellow) and normal estimation results (in green) for a chair model with different noise levels. The normals are rendered as point sprites to indicate their orientation. Downsampled to 50,000 points. Downsampled to 5,000 points.
[…]
Reconstruction and normal estimation results for a chair model with different noise levels
Input pointsGT meshOursiPSRDipoleOursPGROursGCNO
Figure 11. Reconstruction results for four noisy point clouds. Each model is obtained by adding $\sigma=0.5\%$ Gaussian noise to uniform samples. Downsampled to 50,000 points. Downsampled to 5,000 points.
[…]
5.2.2. Thin Structures and Sharp Edges
@@
5.2.3. High-genus Surfaces
We experiment with two high-genus surfaces shown in Fig. 8. Due to their large surface areas, we sample 200,000 for each model to ensure the point clouds can depict the underlying geometric structure. We empirically found that our method and iPSR require more iterations to converge, and thus set the number of iterations for both methods to 300. In Fig. 8 we show the reconstruction results as well as the genus of each mesh. Our method can more faithfully reconstruct topological structures than other methods.
5.2.4. Sparse Points and 3D Sketches
Sparse points can be another challenge due to low sampling rates. Additionally, Huang et al. (2019) studied an extreme case of sparse inputs: 3D sketches, where the points are not only sparse but also highly non-uniform. In Fig. 9 we show four models from Huang et al. (2019). The two uniform sparse point clouds each have 500 points, and the two sketches each have 1,000 points. Our method can obtain plausible normals for both types of point clouds.
Table 4. Quantitative comparisons of estimating normals for noisy models exhibited in Fig. 10 and Fig. 11. Since it does not make sense to talk about “GT normals” of noisy points, we measure only mesh errors. The best results are marked in bold. Downsampled to 50,000 points. Downsampled to 5,000 points.
@@ 5.2.4. Sparse Points and 3D Sketches
Dipole 5.7139 23.4262 6.4231 29.9646 5.1984 32.2373 20.2586 47.9005
@@
Fig. 11 PGR Cup 7.3970 10.3166 Flower 3.2521 4.8484 Buddha 4.4460 18.8665 Vase 3.0096 8.6615
Ours 2.5295 2.2244 1.9771 3.8066 2.3528 8.2716 1.3408 2.3336
@@
Ours 5.0091 5.2786 3.3762 4.8993 3.8669 10.8446 2.6241 2.6057
@@
Figure 12. Reconstruction results of raw range scans from the Stanford scanning repository (Curless and Levoy, 1996; Krishnamurthy and Levoy, 1996; Turk and Levoy, 1994; Gardner et al., 2003) and Lu et al. (2018). Number of points in each model: 2,374,290 (Armadillo); 362,272 (Bunny); 2,109,047 (Dragon); 530,497 (Lady). Please zoom in to see scanner noise patterns and reconstruction details.
[…]
Reconstruction results of raw range scans.
OursiPSRPGRGCNOOursiPSRPGR#Iterations1234510203040
Figure 13. Illustration of the convergence processes of different methods when applied to two uniformly sampled models (3DBenchy and XYZRGB Statuette). We visualize each intermediate result using error maps. Red indicates high error and blue indicates low error. We show the first 40 iterations of each method.
[…]
Error color maps of convergence process.
[…]
Figure 14. Plots of the average angular error and the percentage of correctly oriented normals ($P_{\rm co}$) of the convergence processes of different methods.
[…]
Plots of the average angular error and the percentage of correctly oriented normals ($P_{\rm co}$) of the convergence processes of different method Illustration of the convergence process of different methods.
5.2.5. Noise Resilience
We study the noise resilience of different methods by adding Gaussian noise to clean uniform samples. We use three noise levels $\sigma=0.25\%,0.5\%,1\%$ (w.r.t. the bounding box diagonal) as the standard deviation for Gaussian noise. Following previous settings, each model is sampled with 160,000 points, which are downsampled to 50,000 for PGR (Lin et al., 2022) and 5,000 for GCNO (Xu et al., 2023). We noticed that default settings of several methods do not work well with noisy inputs, and therefore we adjust the parameters to obtain optimal results. For our method, we set $w_{1}=0.02,0.03,0.04$ and $w_{2}=4w_{1}$ for the three noise levels. For iPSR (Hou et al., 2022), we set the point weight to 0.5. For PGR (Lin et al., 2022), we use their recommended settings for noisy data. Note that Dipole (Metzer et al., 2021) and GCNO (Xu et al., 2023) do not have noise-adapting options or parameters, and we keep their default settings. We also change the point weight to 0.5 for the final reconstruction step using SPSR (Kazhdan and Hoppe, 2013).
Fig. 10 shows the reconstructed meshes (in yellow) and the estimated normals (in green) for different noise levels. For dipole propagation (Metzer et al., 2021), even without noise it cannot infer normals with globally consistent orientation, and as noise level increases its normals become completely misoriented. Other methods can infer a roughly correct orientation but are notably worse at preserving the geometric structure, especially at high noise levels. Fig. 11 exhibits four more noisy cases (all models are sampled with noise level $\sigma=0.5\%$). Compared with other methods, ours achieves the best balance between preserving the overall topological structure and surface details. Table 4 reports the quantitative results for the models in Fig. 10 and Fig. 11.
5.2.6. Real-world Scans
Different from synthetic data, real-world scans generally suffer from several kinds of imperfections, such as non-uniform point distribution, scanner noise, outlier points, multi-view misalignment, occluded missing regions, etc. Furthermore, scanner data are usually large in scale, having millions of points. To evaluate the practicability of our method on scanned data, we experiment with three raw range scans (Armadillo, Bunny and Dragon) from the Stanford scanning repository (Curless and Levoy, 1996; Krishnamurthy and Levoy, 1996; Turk and Levoy, 1994; Gardner et al., 2003) and one scan (Lady) from Lu et al. (2018). Note that some raw scans contain background points as well, and we manually set a bounding box to filter out background points.